OpenStudy (anonymous):

What is The closed form solution of the integral of e^(x^2)dx??

OpenStudy (lgbasallote):

closed form solution?

OpenStudy (anonymous):

solution or integral

OpenStudy (lgbasallote):

oh..so many fancy terms

OpenStudy (anonymous):

lol... you are really bouncing around topics, eh? From composition of functions and parabolas yesterday to calculus today? What's up?

OpenStudy (lgbasallote):

a true learner @SmoothMath

OpenStudy (anonymous):

lol online math courses!...killin me

OpenStudy (lgbasallote):

\[\huge \int e^{x^2}dx\] i think i remember this thing as one of those erf function thingies...

OpenStudy (anonymous):

Seriously? Yesterday I taught you how to do f(3)/g(4) and today you're learning integrals? That's ridiculous. I don't believe you.

OpenStudy (anonymous):

Not even simple integrals...

OpenStudy (anonymous):

haha well i dont know what to tell ya..im scholarly...thanks btw igbasallote

OpenStudy (anonymous):

the only thing i know about this problem is that if its an integral problem then i can use the maclaurin expansion for e^u...but thats it hahaha

OpenStudy (anonymous):

use a quadratic equation to find two real numbers that satisfies the situation. The sum of the two numbers is 7 and their product is 14.

OpenStudy (anonymous):

ug this is killing me...there's a reason my username is mathdummy!...

OpenStudy (anonymous):

can anyone confirm that this is the correct indefinite integral?....ex^2 dx=1/4 (square root of pi) erfi(x)??

OpenStudy (anonymous):

What are the bounds of your integral?

OpenStudy (lgbasallote):

so this is killing you...but you arrived with an answer?

OpenStudy (anonymous):

lol ya with some help....its me and 3 other people working on this problem...this is what we collectively got..but im not so sure its right

OpenStudy (anonymous):

The function doesn't have an elementary antiderivative. You can do it numerically for whatever finite bounds you want, but if you're talking about integrating to infinity then it's meaningless because the integral doesn't converge.

OpenStudy (anonymous):

but what about the expansion of the integral at x=infinity? cant that be done?

OpenStudy (anonymous):

No, that function blows up. If you were talking about \[\int e^{-x^2} dx\] then it'd be different, but if the argument of the exponent is +x^2, as x gets larger that thing skyrockets to infinity.

OpenStudy (anonymous):

ok....well thanks for your help!! i'll report back to my team haha

OpenStudy (anonymous):

SmoothMath at once lost for words i see?:)

OpenStudy (anonymous):

What you said earlier is (almost) true, by the way. \[ \int e^{x^2} dx = \frac{1}{2} \sqrt{\pi}\space \text{erfi(x)} \] If that's what you consider closed form, then that's fine....

OpenStudy (anonymous):

sweet!! im surprised we got so close!!! thanks so much for your help!!

OpenStudy (anonymous):

I still think you're trolling me.

OpenStudy (anonymous):

No problem. But just to clarify, that's not any extra insight into the solution, because the whole definition of erfi(x) is \[\text{erfi(x)} = \frac{2}{\sqrt{\pi}} \int_0^x e^{t^2}dt\]

OpenStudy (anonymous):

It's like saying \[x^4 - x^3 + 3x^2 - x =0 \] is solved by \[x = \Omega\] where I define \[\Omega\] to be the solution to the equation \[x^4 -x^3 + 3x^3 -x = 0\]

OpenStudy (anonymous):

well tomorrow im going back to my functions smoothmath haha...this calculus stuff is way to intense for me....and ya jemurray we were just trying to get a base for the indefinte integral....thanks again it REALLY did help

OpenStudy (anonymous):

Who switches each day between learning basic algebra and calculus?

OpenStudy (anonymous):

this girl:) ... i like a challenge but even i have to admit defeat at this point..iv been lost all night